OpenStudy (anonymous):

is this slope correct m= y2 – y1 = 0 -330 = -330 = -3 x2 – x1 110 –0 110

OpenStudy (jdoe0001):

what are the 2 points?

OpenStudy (anonymous):

0,-330 110,0

OpenStudy (jdoe0001):

\(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies \cfrac{0-(-300)}{110-0}\implies \cfrac{300}{110}\implies \cfrac{30}{11}\)

OpenStudy (anonymous):

the equation is m= y2 – y1 = 0 -330 = -330 = -3 x2 – x1 110 –0 110 What is the slope I think it is -3 but not sure.

OpenStudy (anonymous):

it is -330

OpenStudy (jdoe0001):

shoot, I see... then

OpenStudy (jdoe0001):

\(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies \cfrac{0-(-330)}{110-0}\implies \cfrac{330}{110}\implies 3 \)

OpenStudy (anonymous):

thank you. if x=3 than y =1

OpenStudy (jdoe0001):

are you after the equation of the line?

OpenStudy (anonymous):

I need to write from the point slope an equation for a linear inequality

OpenStudy (jdoe0001):

\(\bf slope=m=3\qquad x =3\qquad y =1\\ \quad \\ y-y_1=m(x-x_1)\implies y-(1)=3(x-(3))\implies y-1=3x-9\\ \quad \\ y = 3x-8 \)

OpenStudy (anonymous):

so 3x/-8

OpenStudy (jdoe0001):

no just 3x -8

OpenStudy (anonymous):

how do i substitute x and y in this y=330=? using 3x and -8

OpenStudy (jdoe0001):

not sure what you mean?

OpenStudy (anonymous):

If I need to substitute m for the slope and 330,0 for x and y how do i do it

OpenStudy (anonymous):

are you still helping me?

OpenStudy (jdoe0001):

.... yes..... I'm just unsure on what you're asking :(

OpenStudy (anonymous):

y-330=m(x-0) how would I write the slope (m) using 3x-8 I am not a math pro.

OpenStudy (jdoe0001):

if we disregard the part of x = 3 and y = 1 and instead use the original 2 points for the slope, that is (0,-330) and (110,0) then using the point-slope form it'll be \(\bf slope=m=3\qquad x =0\qquad y =-330\qquad \textit{using (0,-330)}\\ \quad \\ y-y_1=m(x-x_1)\implies y-(-330)=3(x-(0))\implies y+330=3x\\ \quad \\ y = 3x-330\)

OpenStudy (anonymous):

now if I need to write this in an linear inequality using the point slope how do i do it? what steps do I use to come up with linear inequality?

OpenStudy (anonymous):

I need to write an equation for the linear inequality using the point slope. Does this make sense?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

now, notice the line, and let us pick a POINT NOT IN THE LINE let's say the point (100, 0) is not in the line, so let us use that \(\bf point\qquad (100,0)\\ \quad \\ y \quad \square \quad 3x-330\implies 0 \quad \square \quad 3(100)-330\implies 0 \quad \square \quad 300-330\\ \quad \\ 0 \quad \square \quad -30\) so which one do you think is bigger? 0 or -30 ?

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

ahh so we can say that, 0 is GREATER than -30, or \(\bf 0 \quad \square \quad -30\implies 0 \quad > \quad -30\qquad thus\\ \quad \\ \textit{we could also say then that}\qquad y > 3x-330\)

OpenStudy (anonymous):

Thank you I am so confused how to do this problem. I think I am on the right track, however I am probably over thinking this whole algebra thing.

OpenStudy (jdoe0001):

np

OpenStudy (anonymous):

signing off